### v2

# InstructMoLE: Instruction-Guided Mixture of Low-rank Experts for Multi-Conditional Image Generation

Abstract
Parameter-Efficient Fine-Tuning of Diffusion Transformers (DiTs) for diverse, multi-conditional tasks often suffers from task interference when using monolithic adapters like LoRA. The Mixture of Low-rank Experts (MoLE) architecture offers a modular solution, but its potential is usually limited by routing policies that operate at a token level. Such local routing can conflict with the global nature of user instructions, leading to artifacts like spatial fragmentation and semantic drift in complex image generation tasks. To address these limitations, we introduce InstructMoLE, a novel framework that employs an Instruction-Guided Mixture of Low-Rank Experts. Instead of per-token routing, InstructMoLE utilizes a global routing signal, Instruction-Guided Routing (IGR), derived from the user’s comprehensive instruction. This ensures that a single, coherently chosen expert council is applied uniformly across all input tokens, preserving the global semantics and structural integrity of the generation process. To complement this, we introduce an output-space orthogonality loss, which promotes expert functional diversity and mitigates representational collapse. Extensive experiments demonstrate that InstructMoLE significantly outperforms existing LoRA adapters and MoLE variants across challenging multi-conditional generation benchmarks. Our work presents a robust and generalizable framework for instruction-driven fine-tuning of generative models, enabling superior compositional control and fidelity to user intent. Code is available at https://github.com/yanq095/InstructMoLE.
[…]
1 Introduction
The advent of powerful, open-source Diffusion Transformers (DiTs) [peebles2023scalable, esser2024scaling, flux2024] has unlocked unprecedented capabilities in generative AI, fueling a demand for highly specialized and compositional functionalities, from multi-subject composition to personalized content creation [labs2025flux, wu2025qwenimagetechnicalreport, wu2025omnigen2, liu2025step1xeditpracticalframeworkgeneral]. Parameter-Efficient Fine-Tuning (PEFT), particularly Low-Rank Adaptation (LoRA) [hu2022lora], has become the de facto standard for such customization [xiao2024coap]. However, LoRA’s monolithic update structure conflicts with the demands of multi-task fine-tuning, leading to catastrophic forgetting as different task objectives interfere [biderman2024loralearnsforgets, han2024parameterefficientfinetuninglargemodels].
The Mixture-of-Experts (MoE) architecture offers a structured solution to this interference problem. As theoretically grounded by [li2025theory], MoE mitigates catastrophic forgetting by diversifying its experts to specialize in different tasks, which in turn helps to establish, or at least preserve, task-specific expert circuits and balance the loads across them. In the domain of language and multi-modal models, MoE has been extensively studied, leading to established, often task-aware, routing strategies [wumixture, liu2025beyond, li2024mixlora, gou2023mixture, chenoctavius, wu2025routing, sun2025a, fei2024scalingdiffusiontransformers16].
In contrast, its application to diffusion transformers for multi-conditional image generation (e.g. image editing, multi-subject driven generation) remains comparatively underexplored. Initial works like ICEdit [zhang2025context] have adopted conventional token-level routing mechanisms without specializing them for this new problem context. This reliance on token-level routing reveals a potential misalignment with the global nature of image generation instructions. A user’s instruction for compositional generation, for instance, to create a scene with “A baby crawling on the grass, a white horse grazing nearby, and a football helmet” (in Figure 2), establishes a complex set of global semantic relationships. In contrast, token-level routing delegates expert selection to each local image patch independently. This uncoordinated decision-making process can lead to critical failures, often resulting in subjects that are incoherently rendered, spatially fragmented, or have their specified relationships and attributes ignored entirely.
To rectify this misalignment, we propose InstructMoLE, a MoLE framework for multi-conditional image generation built upon a novel, globally consistent routing policy. The core of our framework is Instruction-Guided Routing (IGR), which conditions expert selection entirely on the global semantics of the user’s textual instruction. This mechanism ensures that for any given layer of the model, a single, unified “expert council” is chosen based on the instruction and broadcast to all spatial locations within that layer. This enforces processing consistency at each stage of generation, while critically allowing the model to recruit different sets of specialized experts across different layers, tailoring the computation to the varying levels of abstraction. However, the effectiveness of this globally-applied council hinges on the functional diversity of its constituent experts. To ensure this diversity and prevent representational collapse, we complement IGR with a novel output-space orthogonality loss. This regularizer encourages the learned experts to occupy distinct functional roles, thereby maximizing the compositional power of the selected council. Our main contributions are:
  • We identify a critical challenge in applying MoLE to instruction-based editing and generation: the inherent mismatch between local, token-level routing policies and the global, semantic scope of user instructions. To address this, we propose InstructMoLE, a framework centered on Instruction-Guided Routing (IGR), which aligns expert selection with the holistic intent of the instruction.
  • We introduce a novel output-space orthogonality regularizer to explicitly promote functional diversity among experts. This technique complements standard load-balancing losses by directly penalizing representational redundancy in the expert outputs, thereby mitigating expert collapse and improving the model’s compositional control.
  • We provide extensive empirical validation across multiple challenging benchmarks for multi-conditional image generation. Our results demonstrate that InstructMoLE achieves state-of-the-art performance, particularly in tasks demanding high compositional fidelity and adherence to complex spatial relationships. These findings establish that Instruction-Guided Routing (IGR) is a more effective and principled approach for multi-conditional generation tasks than traditional token-level routing policies.
2 Related Work
2.1 Conditional Generation with Diffusion Transformers.
The advent of Diffusion Transformers (DiTs) has marked a new era for generative modeling, demonstrating remarkable scalability and performance [peebles2023scalable, esser2024scaling]. A significant line of research has focused on enhancing their controllability for complex, instruction-driven tasks. Methods like DreamO [mou2025dreamo] and In-Context Edit [zhang2025context] have enabled sophisticated image customization and editing based on diverse user inputs. Others have introduced lightweight modules for flexible conditioning [zhang2025easycontrol, jiang2025infiniteyou] or tackled data bottlenecks in multi-subject scenarios [wu2025less]. Complementary work has explored compression techniques for attention-based vision models [xiao2023comcat, xiao2025ecospa]. While these works significantly advance the state of conditional generation, they typically rely on a monolithic adaptation strategy, such as a single LoRA [hu2022lora]. This approach struggles when tasked with mastering multiple, potentially conflicting skills simultaneously, creating a clear need for more dynamic, modular adaptation methods.
2.2 Mixture-of-Experts for Parameter-Efficient Fine-Tuning.
The Mixture-of-Experts (MoE) architecture, particularly in its low-rank form (MoLE), offers a powerful solution to the limitations of monolithic fine-tuning. By allocating specialized LoRA experts for different functions, MoLE can mitigate task interference and enhance model capacity efficiently. This paradigm has been extensively and successfully explored in language and multi-modal models for multi-task learning and instruction tuning [dou2024loramoe, gou2023mixture, chenoctavius]. Related efforts have also focused on compressing MoE models through inter-expert pruning and low-rank decomposition [yang2024moe]. Its application to vision, however, is more nascent. While initial works have validated the potential of MoLE for tasks like controllable visual effect generation [mao2025omni], the full power of this architecture is unlocked only through a well-designed routing policy that determines which experts to activate for a given input.
2.3 MoE Design: Routing Policies and Expert Diversity.
The design of an effective MoE system centers on two core challenges: the routing policy and the promotion of expert diversity. The predominant routing approach, inherited from language models, is token-level routing, where each token independently selects experts [shazeer2017outrageously, zhang2025context]. While advanced variants like Expert Choice routing improve dynamics [sunec], they inherit the same fundamental limitation: local, per-token decision-making. This is misaligned with the holistic intent of user instructions in image generation. A second, orthogonal challenge is the common issue of “expert collapse”, where experts become functionally redundant. Standard load-balancing only encourages uniform utilization, not functional diversity. Our work addresses these dual challenges head-on. We introduce Instruction-Guided Routing (IGR), a global policy that aligns expert selection with the user’s instruction, and complement it with an output-space orthogonality loss that directly enforces functional diversity among experts.
3 Methodology
We present InstructMoLE, a Mixture of Low-rank Experts (MoLE) framework for multi-conditional image generation, illustrated in Figure 1. The framework champions a singular, top-down routing principle: a unified Instruction-Guided Routing (IGR) policy is applied consistently across all model layers. This is achieved through a Perceiver-style signal distillation for robust global guidance and an output-space orthogonality loss to maintain functional diversity among experts.
Figure 1: Illustration of the InstructMoLE framework. A global signal, $\mathbf{Z}_{\text{global}}$}, is distilled from the user’s instruction to guide a Router. The Router selects a single, consistent set of LoRA experts, which is then applied to all input tokens.
3.1 InstructMoLE Architecture
Standard LoRA augments a frozen linear layer weight $\mathbf{W}_{0}\in\mathbb{R}^{D_{\text{in}}\times D_{\text{out}}}$ with a single, static low-rank update [hu2022lora, xiao2023haloc]. We replace this static update with a dynamic, sparse mixture of $N$ expert LoRA modules $\{\mathbf{E}_{i}\}_{i=1}^{N}$. Each expert $i$ consists of a down-projection matrix $\mathbf{W}^{i}_{\text{A}}\in\mathbb{R}^{D_{\text{in}}\times r}$ and an up-projection matrix $\mathbf{W}^{i}_{\text{B}}\in\mathbb{R}^{r\times D_{\text{out}}}$, where $r\ll D_{\text{in}},D_{\text{out}}$ is the rank. The function for expert $i$ is defined as $\mathbf{E}_{i}(\mathbf{X})=\mathbf{X}\mathbf{W}^{i}_{\text{A}}\mathbf{W}^{i}_{\text{B}}$. The output of a MoLE layer is therefore:
$\text{MoLE}(\mathbf{X})=\mathbf{X}\mathbf{W}_{0}+\sum_{i=1}^{N}g_{i}(\cdot)\mathbf{E}_{i}(\mathbf{X})$ (1)
where the gating weights $g_{i}(\cdot)$ are determined by a routing network; $\mathbf{X}\in\mathbb{R}^{B\times L\times D_{\text{in}}}$ is the input tensor with batch size $B$ and sequence length $L$.
Token-level vs. Instance-level Routing.
Conventional MoE models employ token-level routing, where each of the $B\times L$ tokens independently selects experts from its local hidden state [fedus2022switch, shazeer2017outrageously, yuanexpert, sunec, zhang2025context], resulting in a dense logits tensor of shape $\mathbb{R}^{B\times L\times N}$. The uncoordinated, per-token decisions create a tension with the inherently global nature of many instructions, leading to critical failures: (i) spatial fragmentation, yielding region-wise style and hue discontinuities; (ii) amplification of high-frequency noise through routing jitter; and (iii) weakened adherence to the user instruction.
To counteract these failure modes, we introduce an instance-level policy that enforces a globally consistent expert assignment at each layer. Instead of fragmented per-token decisions, our router is guided by the global semantics of the user instruction to compute a unified expert council for the entire instance. This council, represented by a compact logits tensor of shape $\mathbb{R}^{B\times 1\times N}$, is then broadcast identically to all $L$ tokens within that layer. We term this mechanism Instruction-Guided Routing. By ensuring that a single, unified expert council is applied to all input tokens within a given layer, IGR directly prevents fragmentation and routing jitter, thereby preserving the semantic and structural integrity of the image.
3.1.1 Instruction-Guided Routing (IGR) Policy
Global Routing Signal ($\mathbf{Z}_{\text{global}}$).
The IGR policy is conditioned on a global signal, $\mathbf{Z}_{\text{global}}$, which is distilled from the editing instruction, $\mathbf{I_{c}}$. To construct a signal that is both semantically robust and compositionally aware, we fuse two distinct text representations. Given an $\mathbf{I_{c}}$, a T5 encoder produces token-level features $\mathbf{H}_{\text{inst}}\in\mathbb{R}^{B\times L_{\text{inst}}\times D_{\text{inst}}}$, where $B$ is the batch size, $L_{\text{inst}}$ is the instruction length, and $D_{\text{inst}}$ is the T5 embedding dimension. Concurrently, a CLIP encoder provides a pooled embedding $\mathrm{CLIP}(\mathbf{I_{c}})\in\mathbb{R}^{B\times D}$. While the pooled embedding offers a strong semantic anchor, it can overlook crucial compositional nuance. For example, in the instruction “Change the woman’s dress to red and the man’s shirt to blue", a pooled embedding might average “red" and “blue", losing the critical association between the colors and their respective subjects. Conversely, token-level features, if naively averaged, risk diluting these key semantics.
We therefore employ a Perceiver-style attentional bottleneck [jaegle2021perceiver] to distill the rich, token-level information from $\mathbf{H}_{\text{inst}}$ into a compact summary before fusing it with the CLIP embedding. Concretely, we first project the T5 tokens into the $D$-dimensional CLIP space using a linear layer $\mathbf{W}_{\text{in}}\in\mathbb{R}^{D_{\text{inst}}\times D}$. A single, learnable latent query, $\mathbf{Q}_{\text{latent}}\in\mathbb{R}^{1\times 1\times D}$, is then used to iteratively query the projected instruction tokens via $S=2$ layers of Perceiver-style attention:
[…]
where $\mathrm{tile}_{B}(\cdot)$ repeats the latent query $B$ times. The final latent, $\mathbf{L}^{(S)}$, now serves as a distilled summary of the compositional details present in the T5 features. This summary is projected with $\mathbf{W}_{\text{out}}\in\mathbb{R}^{D\times D}$, normalized, and then additively fused with the CLIP embedding:
$\mathbf{Z}_{\text{global}}\;=\;\underbrace{\text{LayerNorm}(\mathbf{L}^{(S)}\mathbf{W}_{\text{out}})}_{\text{distilled compositional summary}}\;+\;\underbrace{\mathrm{CLIP}(\mathbf{I_{c}})}_{\text{holistic semantics}}.$ (3)
This design allows the distilled term to contribute fine-grained specificity from the instruction, while the CLIP embedding provides a robust, holistic semantic anchor, together forming a balanced and powerful global routing signal.
IGR Gating Mechanism.
The IGR gating mechanism is instantiated independently at each MoLE layer. We formalize the computation by first describing the process for a single instance $b\in\{1,\dots,B\}$ within a batch, which corresponds to the input tensor slice $\mathbf{X}_{b}\in\mathbb{R}^{L\times D_{\text{in}}}$. The per-instance outputs, $\mathbf{Y}_{b}$, are subsequently stacked to form the full batch output $\mathbf{Y}$.
At a given $l$-th layer, the gating network $\mathcal{G}_{l}:\mathbb{R}^{D}\!\to\!\mathbb{R}^{N}$ generates a logit vector over the $N$ experts from a shared global instruction signal, $\mathbf{Z}_{\text{global},b}$. The Top-$k$ experts are then selected:
$(\mathcal{I}_{b},\mathbf{w}_{b})=\text{Top-}k(\mathrm{Softmax}(\mathcal{G}_{l}(\mathbf{Z}_{\text{global},b})),k).$ (4)
Here, $\mathcal{I}_{b}\subset\{1,\dots,N\}$ is the set of indices for the Top-$k$ experts, and $\mathbf{w}_{b}\in\mathbb{R}^{k}$ contains their associated, unnormalized $\mathrm{Softmax}$ probabilities. $\mathcal{I}_{b}$ and $\mathbf{w}_{b}$ are then broadcast across all $L$ token positions in the sequence. The final output for the $l$-th layer, $\mathbf{Y}_{b}$, is computed by summing the output of a shared linear layer with the weighted sum of the selected expert outputs:
$\mathbf{Y}_{b}=\mathbf{X}_{b}\mathbf{W}_{0}+\sum_{j=1}^{k}w_{b,j}\cdot\mathbf{E}_{\mathcal{I}_{b}[j]}\!\big(\mathbf{X}_{b}\big),$ (5)
where $w_{b,j}$ is the weight for the $j$-th selected expert, whose global index is $\mathcal{I}_{b}[j]$, and $\mathbf{E}_{\mathcal{I}_{b}[j]}$ denotes the corresponding low-rank expert.
3.2 Theoretical Analysis
We provide theoretical insights linking our design choices to established MoE principles.
Proposition 1 (Expert Diversity)
The orthogonality loss $\mathcal{L}_{\text{ortho}}$ enforces a full-rank functional basis over the expert outputs, satisfying the expert diversification requirement established in [li2025theory].
Proof 1
By minimizing $\mathcal{L}_{\text{ortho}}$, the expert output vectors $\{\mathbf{v}_{i}\}_{i=1}^{N}$ are encouraged toward mutual orthogonality. When these vectors are mutually orthogonal, they form a full-rank functional basis in $\mathbb{R}^{B\times L\times D_{\text{out}}}$, ensuring each expert occupies a distinct functional role and reducing the risk of representational collapse.
Proposition 2 (Spatial Consistency)
IGR ensures structural spatial consistency: for any two tokens $x_{i},x_{j}$ in the same instance, $R(x_{i})=R(x_{j})$, where $R(\cdot)$ denotes the routing decision.
Proof 2
IGR computes routing decisions solely from the global signal $\mathbf{Z}_{\text{global}}$, independent of local token states. The routing logits $\mathcal{G}_{l}(\mathbf{Z}_{\text{global},b})$ are broadcast identically to all $L$ tokens, ensuring $R(x_{i})=R(x_{j})$ for all $i,j\in\{1,\dots,L\}$.
These properties directly address the failure modes identified in token-level routing: Proposition 2 eliminates spatial fragmentation, while Proposition 1 ensures the expert pool maintains functional diversity.
3.3 Training Objectives
Our training objective is designed to achieve two goals simultaneously: high-fidelity image generation and the effective, diverse utilization of the expert pool. This is realized through a composite loss function, where the $\lambda$ terms are hyper-parameters:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{flow}}+\lambda_{\text{aux}}\mathcal{L}_{\text{aux}}+\lambda_{\text{ortho}}\mathcal{L}_{\text{ortho}}.$ (6)
Generation Accuracy ($\mathcal{L}_{\text{flow}}$).
The primary objective, $\mathcal{L}_{\text{flow}}$, is the standard flow-matching or diffusion loss of the MM-DiT backbone. This ensures the model accurately synthesizes images that are well aligned with the user’s instruction.
Expert Load Balancing ($\mathcal{L}_{\text{aux}}$).
Following the standard practice for training MoEs [shazeer2017outrageously], we employ an auxiliary load-balancing loss, $\mathcal{L}_{\text{aux}}$, to prevent the model from consistently favoring only a few experts. This loss is a function of two quantities calculated over a batch: $f_{i}$, the fraction of tokens routed to expert $i$, and $p_{i}$, the mean routing probability for that expert:
$\mathcal{L}_{\text{aux}}\;=\;N\sum_{i=1}^{N}f_{i}\cdot p_{i}\,$, (7)
where $N$ is the total number of experts.
Output-space Orthogonality Loss for Functional Diversity ($\mathcal{L}_{\text{ortho}}$).
A critical failure mode in Mixture-of-Experts (MoE) models is expert collapse [fedus2022switch], where distinct experts converge to functionally redundant solutions, thereby nullifying the benefits of the mixture. The standard mitigation, an auxiliary load-balancing loss, only encourages that experts are utilized with similar frequency but provides no explicit mechanism to ensure their functional diversity.
To address this limitation directly, we introduce an orthogonality loss that penalizes the functional similarity between pairs of experts. Our approach operates on the raw, pre-gating outputs. For a given input batch $\mathbf{X}$, we first compute the output of every expert, yielding a set of tensors $\{\mathbf{Y}_{1},\dots,\mathbf{Y}_{N}\}$, where each $\mathbf{Y}_{i}=\mathbf{E}_{i}(\mathbf{X})\in\mathbb{R}^{B\times L\times D_{\text{out}}}$. To measure the functional similarity between any two experts $i$ and $j$, we flatten their respective output tensors into high-dimensional vectors, $\mathbf{v}_{i}=\text{vec}(\mathbf{Y}_{i})$ and $\mathbf{v}_{j}=\text{vec}(\mathbf{Y}_{j})$.
The orthogonality loss, $\mathcal{L}_{\text{ortho}}$, is then defined as the mean of the squared cosine similarities over all unique pairs of these expert output vectors:
$\mathcal{L}_{\text{ortho}}=\frac{1}{N(N-1)}\sum_{i\neq j}\left(\frac{\mathbf{v}_{i}\cdot\mathbf{v}_{j}}{\|\mathbf{v}_{i}\|_{2}\|\mathbf{v}_{j}\|_{2}}\right)^{2}.$ (8)
Minimizing this objective incentivizes the vectors $\{\mathbf{v}_{i}\}_{i=1}^{N}$ to become mutually orthogonal. This, in turn, forces the expert functions $\{\mathbf{E}_{i}\}_{i=1}^{N}$ to learn distinct and complementary representations, effectively preventing expert collapse. In practice, this loss is computed efficiently by calculating the squared off-diagonal elements of the Gram matrix formed by the L2-normalized vectors $\{\mathbf{v}_{i}/\|\mathbf{v}_{i}\|_{2}\}_{i=1}^{N}$. This is crucial for unlocking the full specialization capacity of our MoLE framework.
4 Experiments
4.1 Experimental Setup
Training Data.
Our model’s versatile editing capabilities are a direct result of its comprehensive training mixture. We curate a large-scale dataset by combining publicly available sources with a vast corpus of synthesized data, designed to expose the model to a diverse spectrum of conditional inputs. As illustrated in Figure 6, this includes reference-based tasks (e.g., face swapping, style transfer, re-lighting), multi-subject compositional generation, single-image editing, and spatially controlled generation from both dense (depth, Canny maps) and pose skeleton signals. This diverse training regimen enables a single, unified model to handle a wide array of editing modalities.
Evaluation Benchmarks.
We evaluate all models on a suite of benchmarks, each targeting a distinct capability: OmniContext [wu2025omnigen2] for in-context generation, XVerseBench [chen2025xverse] for multiple subjects-driven generation, GEdit-EN-full [liu2025step1xeditpracticalframeworkgeneral] for single-image editing, and MultiGen-20M [qin2023unicontrol] and COCO Pose 2017 [lin2014microsoft] for spatially controlled generation.
In the following tables, the best results are in bold and the second best are underlined. ($\downarrow$: Lower is better; $\uparrow$: Higher is better).
4.2 Main Results
Figure 2: Qualitative comparison with state-of-the-art models.
Implementation Details.
We fine-tune our model from the Flux.1 Kontext (dev) backbone for 100K steps on 64 NVIDIA H100 GPUs. For the MoLE layers, we use $N=8$ total experts with $k=4$ activated per token and an expert rank of $r=32$. The model is optimized using AdamW with a constant learning rate of $1\times 10^{-4}$ and a global batch size of 256 (4 per device).
Efficiency Analysis.
InstructMoLE achieves computational efficiency through global routing. On H100 GPUs, peak memory usage is 51.25 GB, lower than a token-level routing baseline (Expert Choice, 51.88 GB), as global routing eliminates memory-intensive gather-scatter operations. Inference incurs zero latency overhead compared to the baseline, validating the practical efficiency of our approach.
Qualitative Comparison.
Figure 2 qualitatively compares our model against state-of-the-art methods on several complex editing tasks, visually confirming its superiority. In tasks requiring precise spatial control like Spatial Align and Try-on, competing models often fail to respect geometric constraints, yielding misaligned objects or incorrect poses. In contrast, our model robustly adheres to both dense (depth) and sparse (pose) guidance. Our method also demonstrates a stronger grasp of compositional semantics and identity preservation, particularly in the Multi-Subjects task where it correctly renders all subjects and their relationships. This advantage is most pronounced in the Swap Face task, where InstructMoLE is the only model to produce a coherent result while others fail. This ability to disentangle and execute the spatial, semantic, and identity components of a complex instruction highlights the effectiveness of our approach.
Table 1: Quantitative comparison on the OmniContext benchmark, evaluating both prompt following and identity similarity (ID-Sim), with corresponding qualitative results shown in Figure 3.
Method Prompt Following ID-Sim
Avg.
multi
character
multi
character object
scene
character
scene
character object
single
character
UNO 5.63 16.56 20.65 15.16 16.10 8.49 22.41
DreamO-v1.1 6.10 15.47 19.30 9.46 12.41 8.42 27.77
OmniGen2 7.58 23.81 27.52 18.68 18.20 12.61 42.02
[…]
Figure 3: Qualitative comparison of in-context generation on OmniContext benchmark.
Figure 4: Qualitative comparison of multi-subject driven generation on XVerse benchmark.
Figure 5: Qualitative comparison of single-image editing on GEdit-EN-full benchmark.
In-Context Generation.
As shown in Table 1, InstructMoLE achieves a new state-of-the-art on the OmniContext benchmark, attaining the highest average ID-Sim score. The benchmark evaluates instruction adherence (assessed by GPT-4.1) and identity preservation (ID-Sim) [Deng_2019_CVPR]. While OmniGen2 scores highest in Prompt Following, it suffers from poor identity preservation, scoring 39% lower than InstructMoLE (Ours) on the ID-Sim average. InstructMoLE demonstrates a superior balance, outperforming the strong Flux.1 Kontext baseline on both metrics. Although Flux.1 Kontext scores highest on the multi character sub-metric, this is due to rendering artifacts where subjects are naively concatenated rather than composed into a coherent scene, as shown in Figure 3 and Figure 9.
Multiple Subjects-driven Generation.
We evaluate the model’s ability to compose multiple subjects using XVerseBench [chen2025xverse], a benchmark comprising 20 human identities, 74 unique objects, and 45 different animal species. As detailed in Table 2, we assess performance using three key metrics: instruction adherence (DPG score [hu2024ella]), human identity preservation (Face ID similarity [Deng_2019_CVPR]), and object fidelity (DINOv2 similarity [oquab2024dinov]).
Table 2: Quantitative comparison of multi-subject driven generation. Corresponding qualitative results are presented in Figure 4.
Method DPG ID-Sim IP-Sim Avg.
UNO 74.35 31.96 53.07 53.13
DreamO-v1.1 9.92 8.86 59.93 9.57
OmniGen2 91.91 35.96 0.93 62.93
Flux.1 Kontext (dev) 89.49 53.12 50.78 67.46
InstructMoLE 89.57 60.84 62.81 71.07
InstructMoLE achieves a new state-of-the-art on the benchmark. It outperforms the strong Flux.1 Kontext baseline on the average score by a significant margin of over 3.6. While its instruction adherence (DPG) is on par with other leading methods, our model establishes a decisive advantage in subject fidelity, securing the top scores in both identity preservation (ID-Sim) and object similarity (IP-Sim). More qualitative comparison results are presented in Figure 10.
Single-image Editing.
As shown in Table 3, on the GEdit-EN-full benchmark, which evaluates performance across 11 distinct real-world editing categories by GPT-4.1, InstructMoLE again demonstrates its superior versatility by achieving the highest average score. Our model surpasses the strong Flux.1 Kontext baseline and exhibits a decisive advantage over other leading methods like OmniGen2, with an overall performance improvement of more than 13% over the latter. This strength is particularly evident in its ability to handle both global and local manipulations with high fidelity. It establishes a new state-of-the-art in fundamental tasks such as Color Alter, Material Alter, and Replace, while remaining highly competitive across nearly all other categories.
Generalization to Unseen Tasks.
To assess catastrophic forgetting, we evaluate InstructMoLE on GenEval, a standard text-to-image benchmark. Despite fine-tuning exclusively on editing and spatial tasks, InstructMoLE achieves an overall score of 64.14, outperforming the pre-trained Flux.1 Kontext baseline (63.57). This confirms that the MoLE architecture masters new capabilities without compromising the foundation model’s original generative capacity.
Robustness to Complex Instructions.
On the Complex-Edit benchmark [yang2025textttcomplexeditcotlikeinstructiongeneration], which features highly complex, chain-of-thought instructions, InstructMoLE demonstrates robust generalization. It achieves superior perceptual quality (7.15 vs 7.08) and identity preservation (8.09 vs 8.01) compared to Flux.1 Kontext, confirming that the global routing signal maintains high fidelity even when processing complex and potentially ambiguous instructions.
Table 3: Quantitative comparison of single-image editing. The benchmark comprises 11 fine-grained editing categories reflecting practical user requests, with performance for each assessed by GPT-4.1. The corresponding qualitative results shown in Figure 5.
Model BG Change Color Alt. Mat. Alt. Motion Portrait Style Add Remove Replace Text Tone Avg
[…]
ICEdit 2.73 6.00 4.41 1.74 2.14 5.19 4.41 1.53 4.22 1.58 4.58 3.50
[…]
4.3 Ablation Studies
Implementation Details.
We train all models in our ablation studies under a unified and consistent experimental protocol to ensure a fair comparison, using 8 NVIDIA H100 GPUs for 20K steps.
Evaluation Metrics.
We conduct a comprehensive evaluation across three key domains. For multiple subjects-driven generation and single-image editing, we report scores on the XVerseBench [chen2025xverse], GEdit-EN-full [liu2025step1xeditpracticalframeworkgeneral]. For spatial alignment, we evaluate on 500 samples for each modality, randomly sampled from the evaluation splits of MultiGen-20M [qin2023unicontrol] for depth and Canny edge control, and COCO Pose 2017 [lin2014microsoft] for human pose control. The corresponding metrics are Root Mean Squared Error (RMSE) for depth, F1 score for Canny edges, and an F1 score based on Object Keypoint Similarity (OKS) for pose.
Routing Policy Analysis.
We compare our Instruction-Guided Routing (IGR) against state-of-the-art token-level policies in Table 4, using a high-rank LoRA as a strong baseline under a fair training parameter budget. The results reveal the challenges of token-level routing for this domain; the standard Top-$k$ policy degrades performance, and this limitation is exacerbated by the flexible Expert Race (ER) [sunec], which exhibits severe instability on the Pose F1-task. While the load-balanced Expert Choice (EC) [yuanexpert] proves to be a more stable token-level alternative, it is consistently outperformed by our IGR policy. IGR’s superiority demonstrates that for multi-conditional generation, enforcing a coherent, instance-level signal is more critical than the adaptive, token-wise compute allocation offered by even the strongest token-level methods.
Table 4: Ablation study on routing policies. “Standard” refers to the classic token-level Top-$k$ routing policy [fei2024scalingdiffusiontransformers16]. “Standard+IGR” applies the Standard policy to the early dual-stream blocks and IGR to the later single-stream blocks. “IGR+Standard” applies the reverse configuration.
Model Routing Policy Multi-Subject ($\uparrow$) Single-Subject ($\uparrow$) Depth MSE ($\downarrow$) Canny F1 ($\uparrow$) Pose F1 ($\uparrow$)
Flux.1 Kontext (dev) - 63.87 6.06 63.44 13.72% 0.95%
[…]
MoLE (r=32,N=8,k=4r=32,\,N=8,\,k=4) Standard 64.56 6.07 38.64 28.06% 19.50%
Expert Choice (EC) 65.81 6.07 34.00 38.55% 31.47%
[…]
Standard+IGR 64.86 6.07 36.32 33.33% 9.57%
[…]
Table 5: Ablation study of the IGR signal and the orthogonality loss.
IGR Signal Orthogonality Loss Multi-Subject ($\uparrow$) Single-Subject ($\uparrow$) Depth MSE ($\downarrow$) Canny F1 ($\uparrow$) Pose F1 ($\uparrow$)
$\mathrm{CLIP}(\mathbf{I_{c}})$ w/o 64.54 6.12 35.67 38.12% 36.74%
$\mathrm{CLIP}(\mathbf{I_{c}})$ w/ 64.66 6.12 35.67 40.22% 36.79%
$\mathbf{Z}_{\text{global}}$ w/o 65.65 5.97 33.74 38.92% 37.77%
$\mathbf{Z}_{\text{global}}$ w/ 65.68 6.15 33.34 41.51% 40.97%
Analysis of IGR Components.
We conduct a fine-grained ablation on the core components of our IGR policy in Table 5. The results provide three validations. First, adding the orthogonality loss to the CLIP-only baseline (Canny F1: 38.12% $\rightarrow$ 40.22%) demonstrates its contribution independent of the routing signal. Second, replacing the raw CLIP embedding with our fused signal, $\mathbf{Z}_{\text{global}}$, enhances performance across metrics (Multi-Subject: 64.54 $\rightarrow$ 65.65), confirming the necessity of distilling compositional details from token-level features. Third, combining both components yields the best performance (Canny F1: 41.51%, Pose F1: 40.97%), confirming that signal distillation and orthogonality regularization are critical and complementary.
5 Conclusion
In this work, we addressed multi-task interference in parameter-efficient diffusion models by resolving the conflict between local token-level routing and the global intent of user instructions. We introduced InstructMoLE, a framework built upon Instruction-Guided Routing (IGR), which enforces a globally consistent expert selection within each layer based on the user’s instruction. This mechanism is complemented by an output-space orthogonality loss that promotes expert diversity. Our extensive experiments demonstrate that InstructMoLE significantly outperforms LoRA and other MoLE variants on challenging multi-conditional benchmarks. These results establish that a global, instruction-aware routing policy is a more robust and effective paradigm for complex generative tasks, laying the groundwork for models with more faithful compositional control.